# Mil2: Efficient Cloth Simulation Using Non-distance Barriers and Subspace Reuse

### v3

@@ 8.1. Implementation details
Most parts of the Mil2 framework are matrix-free, except for the subspace solve step of $\mathsf{V}^{\top}(\mathsf{H}+\Delta\mathsf{H})\mathsf{V}$. For solving a linear system of $\mathsf{A}x=b$, the most common practice is to pre-factorize $\mathsf{A}$ and compute $x$ via forward and backward substitutions. In our implementation however, we directly solve $\mathsf{X}\in\mathbb{R}^{r\times r}$ via:
@@
to obtain an approximate of the matrix inverse $\mathsf{X}\approx\mathsf{A}^{-1}/\beta$. The reader should not confuse $\mathsf{A}$ in Eq. (25) with $\mathsf{A}_{i}$ used in local projection (e.g., Eq. (3)). Here, $\mathsf{A}$ refers to the reduced global-stage matrix $\mathsf{A}=\mathsf{V}^{\top}(\mathsf{H}+\Delta\mathsf{H})\mathsf{V}$ with subspace reuse. $\beta$ is a scaling factor estimated as $\beta=\sum|b_{i}|/r$ i.e., the average of absolute values of elements in $b$. Doing so could mitigate numerical drift induced by small or large values in $b$. With $\mathsf{X}$ computed, the system solve becomes a matrix-vector multiplication of $\mathsf{X}b$, which can be parallelized on the GPU. The standard forward/backward substitutions are sequential (even on the GPU). When many LG iterations are needed, computing $\mathsf{X}b$ is more efficient than using the factorized matrix. In our implementation, we simply send $\mathsf{A}$ back to the CPU, compute $\mathsf{X}$, and return it to the GPU. Since the reused subspace has a very low dimensionality, it ensures that the associated computations and CPU-GPU communications are fast and require minimal resources. The whole procedure takes less than $0.1$ ms, which is $30\%-35\%$ faster than factorizing $\mathsf{A}$ on the GPU. Nevertheless, the system solve is not the bottleneck of Mil2. In our implementation, we use rank-2 A-Jacobi as our default GPU solver. A rank-2 A-Jacobi iteration computes two regular Jacobi iterations with one step but using the same computation time (Lan et al., 2022b). Thanks to subspace reuse, weighted SOR is never needed even for stiff simulations (e.g., Fig. 2).
Figure 9. Residual forwarding.  We show an experiment where a piece of tablecloth falls on the teapot. If the TOI computed at the end of a time step is small, the corresponding line search filtering truncates the displacement update for all the vertices, leading to locking/damping artifacts (see highlighted snapshots). Residual forwarding estimates a virtual force to restore the dynamics at non-colliding vertices in the next step. Therefore, the animation looks more natural and realistic.
@@
To fully exploit the capacity of modern GPUs, the broad-phase collision culling leverages a patch-based BVH. Specifically, we build an incomplete BVH whose leaf houses a small patch of the cloth mesh. A patch consists of several inter-connected triangles, normally five to eight. After the initial AABB-based culling at each BVH level, from top to bottom, we exhaustively test triangle pairs between two nearby patches as well as pairs within a patch. At the narrow phase stage, if a full CCD is needed we solve $t_{\text{TOI}}$ for each primitive pair using the polynomial solver proposed in (Yuksel, 2022). When partial CCD suffices, we compute inner products of Eq. (10) at pre-selected sample points plus the projection points at $t=0$. We periodically check if a denser sampling is needed given the current system velocity (since the time step size is assumed fixed). Partial CCD is more efficient for cloth models of higher resolutions. We noted that, as long as the total number of DOFs exceeds $100$K, very few sample points (e.g., three) work well for partial CCD in most simulation scenarios.
8.2. EXP vs DBB
The first test we would like to show is a comparison between the distance-based barrier (DBB) and non-distance barrier i.e., EXP. The major difference lies in the fact that EXP allows the stiffness of the collision constraint to be timely adjusted during LG iterations based on inexpensive partial CCD. The snapshots of the resulting simulation are reported in Fig. 8. We also compare the total number of LG iterations using EXP and DBB for this twisting test, and the plots are shown in Fig. 10. It can be seen that EXP helps reduce iterations, as each collision constraint is more likely to find an appropriate weight during iterations. EXP is not sensitive to the sample density ($\rho$). To this end, Fig. 10 also provides iteration counts for EXP with different sampling densities. We can see that EXP still works well even using one sample point at the center of $\Omega_{\lambda}$.
@@ 8.2. EXP vs DBB
Figure 10. DBB vs EXP. We plot the total number of LG iterations at each time step when twisting the cloth (as in Fig. 8) using DBB and EXP. It can be seen that EXP uses $25\%$ to $30\%$ fewer iterations on average than DBB. Partial CCD is not sensitive to the sample density. As shown in the figure, increasing or decreasing partial CCD samples do not vastly alter the convergence behavior of the simulation. Different sample patterns are also visualized in the figure.
@@ 8.3. Residual forwarding
Mil2 exploits residual forwarding to alleviate the damping and locking artifacts induced by line search filtering under a small TOI. Fig. 9 shows a side-by-side comparison with and without using RF. In this experiment, a piece of tablecloth falls on a teapot. The cloth has $16,641$ vertices and $32,768$ triangles. We use a small-size line search filtering at the end of each time step by setting $\alpha=0.1$. The resulting animation suffers from severe damping artifacts since such a close-to-zero TOI freezes the deformation of all the triangles. This observation is highlighted in the figure. Residual forwarding recovers most missing dynamics and produces plausible results. Please refer to the supplementary video for an animated comparison.
@@
Figure 11. “Animal crossing”.  The subspace reuse technique allows Mil2 to handle stiff materials and deformable bodies with ease. This example includes $33$ animal toys, 10 elastic ribbons, $678$K DOFs, and one million elements. Both Mil2 and PD-IPC (Lan et al., 2022b) produce penetration-free animations, while Mil2 is 12×12$\times$12 × faster than PD-IPC.
8.4. Mil2 vs PD-IPC
@@
Lan et al. (2022b) also combine PD with incremental potential contact (PD-IPC) for deformable and cloth simulation. Compared with Mil2, PD-IPC is based on DBB, and a full CCD must be used every time the constraint set is to be updated. PD-IPC solely relies on A-Jacobi for the global solve. For stiff simulations, one must use weighted A-Jacobi to ensure the iteration does not diverge. This is not a problem for Mil2 since the subspace solve removes most low-frequency errors (e.g., as shown in Figs. 2 and 5). As a result, Mil2 outperforms PD-IPC by a significant margin in general. Fig. 11 shows a deformable body simulation result. In this example, we have 33 deformable animal toys, 10 elastic ribbons, and over one million elements, falling into a glass tank. Some animal toys are five times stiffer than others. Such heterogeneous materials/bodies are particularly challenging for PD-IPC as the weight of the A-Jacobi must be set conservatively ($\omega\geq 0.85$), which impairs the convergence. In this example, Mil2 is 12×12$\times$12 × faster than PD-IPC.
@@
Figure 13. Funnel.  We place a rigid and heavy ball (simulated using affine body dynamics (Lan et al., 2022a)) on a funnel covered by three layers of cloth. The cloths hold the sphere under high frictional contacts (top). When the friction is not strong enough, all the cloths fall on the ground eventually (bottom). The simulation involves nearly $700$K DOFs, and Mil2 runs at $6.6$ FPS.
8.5. Mil2 vs BFGS-PD
Another relevant competitor is from a recent contribution by Li et al. (2023) or BFGS-PD. BFGS-PD combines Jacobi iteration with BFGS for global matrix update in a spline-based subspace. BFGS-PD is also based on DBB and uses reduced IPC Hessian for the global-stage solve. BFGS requires a low-rank matrix update at LG iterations, and the B-spline subspace is of high dimension i.e., $r\approx 10,000$. As a result, BFGS-PD is much slower than the pre-computed subspace update used in Mil2. A concrete experiment is reported in Fig. 12, where we pull two intertwining cloth strips to make a tight knot. There are $104$K vertices on the strips. Mil2 runs this example at an interactive rate of $11.8$ FPS, while BFGS-PD needs several seconds for each frame.
Mil2 is able to capture frictional contacts/collisions per user specifications. We follow the strategy used in (Li et al., 2023) to leverage barrier Hessian as the damping matrix to simulate frictional surfaces with different friction coefficients. It is also convenient to model friction as a tangential velocity damping. Fig. 13 reports a simulation under different frictional setups. Three pieces of cloth cover a funnel, and we place a rigid/heavy ball on the funnel. We use affine body dynamics (Lan et al., 2022a) to simulate the ball’s motion. There are nearly $700$K DOFs in this scene, and Mil2 produces physically correct dynamics under different friction settings. The simulation runs at $6.6$ FPS.
Figure 14. Stack on the teapot.  We stack ten pieces of tablecloth on the teapot simultaneously. The free fall of stacking clothes generates a large number of inter-tangled cloth-cloth collisions. Mil2 robustly handles this challenging scene and produces high-quality animation.
[…]
Figure 16. Barbarian ship in cloth.  A barbarian ship falls on three thin rods and slides down to a cloth. The ship model has complex geometries with a lot of local details. They not only generate intensive self-collisions but also lead to high-frequency cloth deformation when the ship reaches the cloth. Mil2 robustly handles this case, and the resulting simulation is $12$ FPS.
Figure 17. Cloth blender.  We drop 14 clothes into a bowl-like container. The blender then rotates and stirs all the clothes. This example generates fast collisions between clothes and the collider, as well as a large number of cloth-cloth collisions. The simulation has $547$K DOFs and runs at about $9$ FPS.
8.6. More experiments
We have tested Mil2 in a wide range of simulation scenes. The detailed time statistics of all the experiments shown in the paper are reported in Tab. 1. Figs. 14, 15, and 17 show three tests with massive collisions. In Fig. 14, we test the robustness of Mil2 under extensive stacking. In this test, ten falling tablecloths hang on the teapot. This simulation generates a lot of interleaving collisions when clothes stack under gravity. Fig. 15 reports another test where we fold four clothes by vertically dropping them on the desk. As the cloth makes contact with the desk surface, it folds in a zig-zag pattern, resulting in numerous self-collisions, particularly edge-edge collisions. Fig. 17 mimics a blender with 14 clothes in a bowl-shaped container. The scene also involves a large number of dynamic collisions. Mil2 produces interesting animations in all of those experiments, and the results are free of any inter-penetration.
While Mil2 is designed for cloth simulation, it is also capable of simulating deformable objects. In addition to the example reported in Fig. 11, we give another experiment with two-way coupling between cloth and deformable body under complex collisions. As shown in Fig. 16, a barbarian ship slides down over three thin rods and is held by a cloth on the right. The complex geometry of the ship, e.g., the masts, canvases, and paddles at the sides, induces fine-grain deformations, which are two-way coupled with the cloth. Our method robustly captures those effects, and the simulation is interactive at $12$ FPS.
Cloth animations play a pivotal role in the realm of digital fashion and design. Mil2 enhances this aspect by enabling virtual characters to interact seamlessly with a variety of garments, yielding high-quality simulations. In demonstration of its capabilities, we present two additional examples. The first features a character executing a kicking action, illustrating the dynamic interaction between the motion and the garment. The second example showcases a virtual model on a fashion runway. This model, attired in a knee-length skirt, walks to the front of the stage before turning around to return. Throughout this sequence, Mil2 meticulously captures the intricacies of the motion.
@@ 8.6. More experiments
Table 1. Time performance.  We report detailed time statistics for experiments mentioned in the paper. ##\## Bdy gives the total number of objects in the scene. ##\## DOF is the total number of simulation DOFs. ##\## Ele. records the total number of elements (i.e., triangles and tetrahedrons). ##\## Con. and ##\## Col. are the numbers of elasticity constraints ($\Psi$) and collision constraints ($B$) on average during the simulation. In the column of ##\## Col., the first quantity reports the number of collision constraints involved in CCD and the second number is the total number of primitive pairs after the broad phase collision detection. $\bm{\bar{r}|r}$ represent the subspace size and reused subspace size. $\bm{\Delta t}$ is the time step size. Pre. is the pre-computation time (measured in seconds). Broad gives the time used for the broad phase collision detection using patch-based BVH. Narrow is the total time used for the narrow phase CCD, which includes partial CCD (P-CCD) and full CCD (F-CCD). LG (##\##) gives the total time (first row) used and total number of LG iterations (second row) on average for each step. Misc. corresponds to other additional computations. Except for Pre., other timings are measured in milliseconds. FPS is the overall FPS of the simulation.
Scene ##\## Bdy ##\## DOF ##\## Ele ##\## Con. ##\## Col. $\bm{\bar{r}|r}$ $\bm{\Delta t}$ Pre. Broad Narrow P-CCD F-CCD LG (##\##) Misc. FPS
@@
$9.7$ $8.9$
@@
$10.1$ $5.4$

### v4

@@ 8.1. Implementation details
Most parts of our framework are matrix-free, except for the subspace solve step of $\mathsf{V}^{\top}(\mathsf{H}+\Delta\mathsf{H})\mathsf{V}$. For solving a linear system of $\mathsf{A}x=b$, the common practice is to pre-factorize $\mathsf{A}$ and compute $x$ via forward and backward substitutions. In our implementation however, we directly obtain $\mathsf{X}\in\mathbb{R}^{r\times r}$ via:
@@
to get an approximation of $\mathsf{X}\approx\mathsf{A}^{-1}/\beta$. The reader should not confuse $\mathsf{A}$ in Eq. (25) with $\mathsf{A}_{i}$ used in local projection (e.g., Eq. (3)). Here, $\mathsf{A}$ refers to the reduced global-stage matrix $\mathsf{V}^{\top}(\mathsf{H}+\Delta\mathsf{H})\mathsf{V}$ with subspace reuse. $\beta$ is a scaling factor estimated as $\beta=\sum|b_{i}|/r$ i.e., the average of absolute values of elements in $b$. Doing so could mitigate numerical drift induced by small or large values in $b$. This is because calculating $\mathsf{X}$ via $\mathsf{AX}=\mathsf{Id}$ implicitly assumes the r.h.s. of the system is around one.
@@
With $\mathsf{X}$ computed, the system solve becomes a matrix-vector multiplication of $\mathsf{X}b$, which can be parallelized on the GPU. The standard forward/backward substitutions are sequential (even on the GPU). When many LG iterations are needed, computing $\mathsf{X}b$ is more efficient than using the factorized matrix. In our implementation, we simply send $\mathsf{A}$ back to the CPU, compute $\mathsf{X}$, and return it to the GPU. Since the reused subspace has a very low dimensionality, it ensures that the associated computations and CPU-GPU communications are fast and require minimal resources. The whole procedure takes less than $0.1$ ms, which is $30\%-35\%$ faster than factorizing $\mathsf{A}$ on the GPU. Nevertheless, the system solve is not the bottleneck. In our implementation, we use rank-2 A-Jacobi method, which computes two regular Jacobi iterations with one step but using the same computation time (Lan et al., 2022b). Thanks to subspace reuse, weighted SOR is never needed even for stiff simulations (e.g., Fig. 2). The base of exponential NDB (i.e., Eq. (7)) is set as $K=2$ in our experiments. Setting $K$ to $3$ seems to produce a similar result. However, an over-aggressive $K$ could negatively impact the convergence. In this case, this exponential NDB behaviors like a geometric projection.
@@
To fully exploit the capacity of modern GPUs, the broad-phase collision culling leverages a patch-based BVH. Specifically, we build an incomplete BVH whose leaf houses a small patch of the cloth mesh. A patch consists of several inter-connected triangles, normally five to eight. After the initial AABB-based culling at each BVH level, from top to bottom, we exhaustively test triangle pairs between two nearby patches as well as pairs within a patch. At the narrow phase stage, if a full CCD is needed we solve $t_{TOI}$ for each primitive pair using the polynomial solver proposed in (Yuksel, 2022). For partial CCD, we compute inner products of Eq. (10) at pre-selected sample points plus the projection points at $t=0$. We periodically check if a denser sampling is needed given the current system velocity (since the time step size is assumed fixed). Partial CCD is more efficient for cloth models of higher resolutions. We noted that, as long as the total number of DOFs exceeds $50$K, very few sample points (e.g., three) work well for partial CCD in most simulation scenarios.
@@
Figure 9. NDB vs DBB. We plot the total number of LG iterations at each time step when twisting the cloth (as in Fig. 8) using DBB and exponential NDB. It can be seen that our NDB formulation uses $25\%$ to $30\%$ fewer iterations on average than DBB when allows the constraints to be adaptively re-weighted in time. Partial CCD is not sensitive to the sample density. As shown in the figure, increasing or decreasing partial CCD samples do not vastly alter the convergence behavior of the simulation. Different sample patterns are also visualized in the figure.
8.2. NDB & DBB
The first test we would like to show is a comparison between the distance-based barrier (DBB) and exponential non-distance barrier i.e.,NDB. The major difference lies in the fact that NDB allows the weight of the collision constraint to be timely adjusted during LG iterations based on inexpensive partial CCD. The snapshots of the resulting simulation are reported in Fig. 8. We also compare the total number of LG iterations using NDB and DBB for this twisting test, and the plots are shown in Fig. 9. It can be seen that this adaptive weighting strategy helps reduce iterations, as each collision constraint is more likely to find an appropriate weight during iterations. NDB strategyis not sensitive to the sample density ($\rho$). To this end, Fig. 9 also plots iteration counts for NDB with different sampling densities. We can see that NDB works well even using one sample point at the center of $\Omega_{\lambda}$.
Figure 10. Residual forwarding.  We show an experiment where a piece of tablecloth falls on the teapot. Residual forwarding estimates a virtual force to restore the dynamics at non-colliding vertices in the next step. It produces reasonably good results when the residual errors are moderate. RF fails to handle all the errors with a highly constrained iteration cap. As we lower the per-step iteration count, we observe more artifacts even using RF.
@@ 8.3. Residual forwarding
Our method employs RF to mitigate damping and locking artifacts caused by limited time budgets. As discussed in Sec. 6, RF can enhance animation quality to a certain extent by “inheriting” unresolved residual errors from one time step to the next. The effectiveness of RF is primarily up to the magnitude of these errors. We observe that small and localized errors, resulting from early termination or line search filtering under a small TOI, are generally well-managed by RF. Another factor is the capability of minimizing residuals at the current time step. If the solver does not fully converge at the current step, the carried-over residuals may exacerbate simulation inaccuracies. RF is particularly designed for applications facing strict time constraints, such as gaming, where high-speed collisions might cause spiky FPS drops.
To better illustrate the effectiveness of RF, we simulate a scene where the cloth drops on a teapot (Fig. 10). There are $120$K DOFs on the cloth. After the cloth comes in contact with the spout of the teapot, the inertia effect further moves the cloth towards the right, and the cloth eventually settles at the recess between the spout and the body of the teapot. When we have the time budget to complete 50 iterations per time step (which do not fully converge the solver), RF produces a high-quality result nearly identical to the ground truth – the one generated using a fully converged Newton IPC solver. Our method runs at $20$ FPS, which is 500×500$\times$500 × faster than (Li et al., 2021a). If we half the iteration number and exit the current time step with the line search filtering, the simulator is unable to deal with all the residuals, and we can see dampened cloth movement after it touches the teapot. Because of the early termination, the cloth “sticks” to the tip of the spout as the rest part of the cloth is locked (see highlighted area in the figure). Lastly, we further lower the iteration cap to only five. We observe more severe damping artifacts. Because of the subspace resue, our method still handles low-frequency residuals well, but most high-frequency information is lost due to an insufficient number of A-Jacobi iterations. The entire cloth exhibits rubber-like dynamics, and the collision is highly inelastic. This comparison is available in the supplementary video.
@@
Figure 11. “Animal crossing”.  The subspace reuse technique allows our solver to handle stiff materials and deformable bodies with ease. This example includes $33$ animal toys, 10 elastic ribbons, $678$K DOFs, and one million elements. Both our method and PD-IPC (Lan et al., 2022b) produce penetration-free animations, while our method is 12×12$\times$12 × faster than PD-IPC on RTX 3090.
8.4. Comparison with existing methods
PD is a popular framework, based on which many excellent cloth and deformable simulation algorithms have been developed. In this subsection, we compare our method with several representative PD-based cloth simulation methods, including PD-IPC (Lan et al., 2022b), PD-BFGS (Li et al., 2023), and PD-Coulomb (Ly et al., 2020).
Our method vs. PD-IPC
@@ 8.4. Comparison with existing methods
PD-IPC (Lan et al., 2022b) is a full GPU simulator. It also integrates PD with IPC for deformable and cloth simulation. Unlike our method, PD-IPC uses DBB, and a full CCD must be invoked every time the constraint set is to be updated. PD-IPC solves the global step system only using A-Jacobi iterations on the GPU. For stiff simulations, one must adopt a close-to-one SOR weight to ensure A-Jacobi iterations do not diverge. This is not an issue for us since the subspace solve removes dominant low-frequency errors. As a result, our method outperforms PD-IPC by a significant margin in general. Fig. 11 reports a deformable body simulation result. In this example, we have 33 deformable animal toys, 10 elastic ribbons, and over one million elements falling into a glass tank. Some animal toys are five times stiffer than others. Such heterogeneous materials/bodies are particularly challenging for PD-IPC as the SOR weight of the A-Jacobi must be set conservatively ($\omega\geq 0.85$), which impairs the convergence. In this example, our method is 12×12$\times$12 × faster. As NDB does not depend on the distance, updating NDB can be processed using partial CCD. This strategy significantly speeds up the evaluation of barrier weighting and reduces more than 60% of computation time.
@@
Table 1. Experiment statistics.  We report detailed time statistics for experiments mentioned in the paper. ##\## B gives the total number of objects in the scene. ##\## DOF is the total number of simulation DOFs. ##\## Ele. records the total number of elements (i.e., triangles and tetrahedrons). ##\## Con. and ##\## Col. are average numbers of elasticity constraints and collision constraints during the simulation. In the column of ##\## Col., the first quantity reports the number of collision constraints involved in CCD and the second number is the total number of primitive pairs after the broad phase collision detection. $\bm{\bar{r}|r}$ represent the subspace size and reused subspace size. $\bm{\Delta t}$ is the time step size. Pre. is the pre-computation time (measured in seconds). The column B—N—P—F gives the timing information (in milliseconds) used for collision detection. Specifically, B is the broad phase time, and N is the narrow phase time. The narrow phase also includes partial CCD (P) and full CCD (F) procedures. $\rho_{m}$ gives the mass density of the cloth (and the deformable objects e.g., in Figs. 11 and 17). The unit is $\frac{kg}{m^{2}}$ for the cloth (and $\frac{kg}{m^{3}}$ for deformable objects). $\kappa$ (in $MPa$) gives the stretching and bending stiffness of the cloth (and Young’s modulus of deformable objects). $\|\Delta x\|$ is the convergence condition. LG (##\##) gives the total time (first row) used and total number of LG iterations (second row) on average for each step. Misc. corresponds to other additional computations. Except for Pre., other timings are measured in milliseconds. FPS is the overall FPS of the simulation.
Scene ##\## B ##\## DOF ##\## Ele ##\## Con. ##\## Col. $\bm{\bar{r}|r}$ $\bm{\Delta t}$ Pre. B—N—P—F $\rho_{m}$ $\kappa$ $\|\Delta x\|$ LG (##\##) Misc. FPS
@@
$10.2$ $5.2$
Cover the ship
(Fig. 17)
[…]
$120|30$ $\frac{1}{150}$ $31.8$ $8|8.3|1.2|7.1$ $0.8|0.9$
$160|8\cdot 10^{3}$
$2\cdot 10^{-4}$
$5\cdot 10^{-4}$
$117.8$
$(51)$
$5.3$ $7.2$
Funnel
(Fig. 18)
[…]
$120|30$ $\frac{1}{150}$ $32.8$ $17|13|2|9$ $0.3$
$160$
$3\cdot 10^{-4}$
$5\cdot 10^{-4}$
$112.4$
$(52)$
$8.5$ $6.6$
@@
$12.4$ $4.8$
Our method vs PD-BFGS
Another relevant competitor is from a recent contribution by Li et al. (2023) or PD-BFGS. PD-BFGS designs a two-step global solve combining Jacobi iteration with BFGS method in a spline-based subspace. PD-BFGS is also based on DBB, and thus slower than our method for IPC-based collision processing. At the global stage, PD-BFGS employs a reduced quasi-Newton procedure in a B-spline subspace. This subspace is of high dimension i.e., $r\approx 10,000$. As a result, PD-BFGS is much slower than the pre-computed subspace update used in our simulation. A concrete experiment is reported in Fig. 12, where we pull two intertwining cloth strips to make a tight knot. There are $104$K vertices on the strips. Our method handles this example at an interactive rate of $11.8$ FPS, while PD-BFGS needs seconds for each frame.
Figure 13. Drape on a sphere.  A tablecloth of $40$K vertices drapes on the sphere. We compare our method with PD-Coulomb (Ly et al., 2020). PD-Coulomb models Coulomb friction, while our method uses a lagged friction Hessian at each LG iteration as in (Li et al., 2023).
Figure 14. Stack on the teapot.  We stack ten pieces of tablecloth on the teapot simultaneously. The free fall of stacking clothes generates a large number of intertwining cloth-cloth collisions. EXP and partial CCD robustly handle this challenging scene and produce high-quality animation.
Figure 15. Just folding.  In this example, four clothes fall vertically on the desk one by one with different orientations. The contact between the cloth and the desktop folds the cloth with complex self-collisions. This is a good stress test because the velocity of falling cloth is high. The resulting multi-layer self-collisions are particularly challenging. There are nearly $500$K DOFs involved, and the simulation runs at $12$ FPS.
Our method vs PD-Coulomb
Our method is compatible with various friction models. In our implementation, we follow the strategy used in (Li et al., 2023), which leverages a quadratic barrier proxy to estimate the Hessian of the friction energy to simulate frictional surfaces with different friction coefficients – in a way similar to the original IPC. Alternatively, Ly et al. (2020) (PD-Coulomb) show that it is possible to incorporate full Coulomb friction by a novel splitting strategy so that the global matrix is kept constant. To this end, we show a side-by-side comparison between our method and PD-Coulomb (Ly et al., 2020).
Figure 16. Cloth blender.  We drop six clothes into a bowl-like container. The blender rotates and stirs all the clothes. This example generates fast rotational collisions between clothes and the collider, as well as a large number of cloth-cloth collisions. The simulation has $820$K DOFs and runs at about $5$ FPS.
Figure 17. Cover the ship.  We cover a deformable barbarian ship with a piece of heavy cloth ($\rho_{m}=0.8kg/m^{2}$). The soft masts and ladders on the ship bend under the weight of the cloth cover. The ship-cloth contacts generate fine wrinkles that delineate the shape features of the ship. There are $540$K DOFs on the cloth and $390$K DOFs on the barbarian ship. With $\Delta t=1/150$, our method resolves all the cloth-cloth, cloth-ship, and ship-ship collisions and runs at $7.2$ FPS.
In this experiment, we drape a piece of square cloth of $40$K vertices on a sphere. The size of the cloth is $1m\times 1m$, and the time step is $\Delta t=1/200$ sec. It can be seen from Fig. 13 (and also see the supplementary video) that both methods capture the frictional contact between the cloth and the sphere and produce similar animation results. In fact, the splitting method proposed in (Ly et al., 2020) can also be integrated with our pipeline. PD-Coulomb adopts DCD, and we follow the default setting as in the published code of (Ly et al., 2020). The collision tolerance is set as $1mm$. This means as long as the distance between two triangles is smaller than $1mm$, DCD will generate a contact. While this is a conservative configuration, considering the movement of the cloth is moderate, one can still spot minor inter-penetration among the cloth triangles. On the other hand, our method uses CCD and a line search filtering at the end of each step, which guarantees all the triangles are separate.
It should be mentioned that PD-Coulomb is CPU-based, using OpenMP to speed up pre-factorized global solve, while our method is fully GPU-based. As a result, our method runs faster (at $25$ FPS) than PD-Coulomb by 30 times. Such performance difference largely comes from the hardware platform. Nevertheless, our method is orthogonal to PD-Coulomb as we do not focus on friction modeling but more on solver optimization and collision detection.
Figure 18. Funnel.  We place a rigid and heavy ball (with affine body dynamics (Lan et al., 2022a)) on a funnel covered by three layers of cloth. The cloths hold the sphere under high frictional contacts (top). When the friction is not strong enough, all the cloths eventually fall on the ground (bottom). The simulation involves nearly $700$K DOFs, and our method runs at $6.6$ FPS.
8.5. More experiments
We have tested our method in a wide range of simulation scenes. The detailed time statistics of all the experiments shown in the paper are reported in Tab. 1. Figs. 14, 15, and 16 show three tests with massive collisions. In Fig. 14, we test the robustness of our method under extensive stacking. In this test, ten falling tablecloths hang on the teapot. This simulation generates a lot of overlapping collisions when clothes stack under gravity. Fig. 15 reports another test where we fold four clothes by vertically dropping them on the desk. As the cloth makes contact with the desktop, it folds in a zig-zag pattern, resulting in a large number of self-collisions, particularly edge-edge collisions. This experiment is a good stress test showing the robustness of a cloth simulator. Fig. 16 mimics a washer with six clothes in a bowl-shaped container. The scene also involves a large number of dynamic collisions. Our method produces interesting animations in all of those experiments, and the results are free of any inter-penetration.
Our method is also capable of simulating deformable objects. In addition to the example reported in Fig. 11, we give another experiment with two-way coupling between cloth and deformable body under complex collisions. As shown in Fig. 17, a heavy cloth falls onto a barbarian ship. The ship base is fixed, and it has soft masts and ladders. As the cloth falls down, we can see detailed wrinkles at the contacting area between the cloth and the ship. There are $540$K DOFs on the cloth and $390$K DOFs on the ship. Our method generates interesting animation results at 7.2 FPS.
Fig. 18 shows two more simulations under different frictional setups. Three pieces of cloth cover a funnel, and we place a rigid/heavy ball on the funnel. We use affine body dynamics (Lan et al., 2022a) to simulate the ball’s motion. There are nearly $700$K DOFs in this scene, and our method produces plausible dynamics under different friction settings. The simulation runs at $6.6$ FPS. As discussed, we can switch to a more accurate frictional model either as in (Ly et al., 2020), which is based on complementarity programming or as in (Li et al., 2020), which is based on the interior-point method.
Cloth animations play a pivotal role in the realm of digital fashion and design. The proposed method enhances this aspect by enabling virtual characters to interact seamlessly with a variety of garments, yielding high-quality simulations. As a demonstration of its capabilities, we report two additional examples. The first features a character executing a kicking action, illustrating the dynamic interaction between the motion and the garment (Fig. 19). This is a numerically challenging test as the body undergoes swift movements. Most DCD-based methods are unable to capture such high-velocity garment dynamics. The second example showcases a virtual model on a fashion runway as shown in Fig. 1. This model, attired in a knee-length skirt, walks to the front of the stage before turning around to return. Throughout this sequence, our method meticulously captures the intricacies of the motion. Please refer to the supplementary for the animated results, where we report two sequences of animations